# Iterative Refinement of Project-Level Code Context for Precise Code Generation with Compiler Feedback

### v2

@@ Appendix A More details about the Error Cases
Table A1: A complete list of frequently occurred errors reported by the compiler
@@
Appendix B Errors Reported by the Compiler
@@
Each error is identified by an error code and described with an error message. The error code is a numerical identifier that reflects the error’s nature, while the error message provides a concise description of the error. Pylint recognizes 133 distinct error types, and only 12 of these are reported frequently on the CoderEval dataset. We have categorized these errors into four groups based on their characteristics, detailed in Table 1. The FUNC error category is also included in the error distribution analysis but is excluded from the compiler feedback pipeline. This exclusion is because such errors can only be identified with execution, a process that extends beyond the scope of static analysis and is unsuitable for the real-time generation and refinement pipeline. The specific types of each error and corresponding error codes are documented in Table A1.
Table A2: Tables pre-computed by ProCoder for error correction.
@@
Table A3: A complete list of demonstration examples prompted to the language model
@@
Appendix C Technical Details of Structural Query System
C.1 The Design Principle
In ProCoder, we employ the Structural Query Language (SQL) to perform structured queries on code repositories. The SQL syntax is derived from CodeQL, which is an analysis tool designed for vulnerability mining through structured queries, capable of conducting complex analyses on code repositories. We extracted 11 error codes from compiler (shown in Table A1), each corresponding to one of four error categories as in Section 2.2. Two authors manually inspected these error codes, and constructed 6 SQL data tables from the project graph, based on the characteristics of the errors. All tables are presented in Table  A2. Whenever an error is reported by the compiler, the LLM generates an SQL query statement for it, where the modules and functions to be queried are extracted using a tree-sitter-based parser  [^1] https://github.com/tree-sitter/tree-sitter-ql. In the ProCoder with iteration rounds set to 3, 93.2% of the SQL queries generated by the LLM pass the parsing utility, and these queries are similar to the demonstration example SQL queries. Subsequently, according to the entity from the query, it is mapped to one of the six SQL data tables. We store these six tables in a relational database for flexible retrieval.
C.2 Demonstration Examples of Structural Queries
We present a demonstration example in Section 3.3 to compose the structural query. The example focuses on identifying and addressing instances of missing or incorrectly utilized context entries. We detail several example error messages along with their corresponding structural queries. A comprehensive list of these examples can be found in Table A3.
Appendix D More Analytic Experiments
D.1 Ablation on Components
Table A4: Pass rates of ProCoder on Class Runnable Test.
Data Split Class Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 28.00 44.92 49.09
ProCoder w/o CF and SQL (RepoCoder) 35.45 40.46 41.82
w/ CF 30.00 42.58 45.45
w/ CF+SQL 30.36 44.61 49.09
w/ CF+Semantic 31.45 45.00 47.27
Table A5: Pass rates of ProCoder on File Runnable Test.
Data Split File Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 30.29 43.58 47.06
ProCoder w/o CF and SQL (RepoCoder) 29.41 34.61 36.76
w/ CF 26.03 37.41 42.65
w/ CF+SQL 26.76 37.68 39.71
w/ CF+Semantic 27.35 39.90 45.45
Table A6: Pass rates of ProCoder on Project Runnable Test.
Data Split Project Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 21.30 36.73 39.13
ProCoder w/o CF and SQL (RepoCoder) 16.96 19.57 21.74
w/ CF 14.35 25.62 30.43
w/ CF+SQL 15.65 25.12 30.43
w/ CF+Semantic 17.83 28.50 34.78
To explore the impact of each component within ProCoder on overall performance, we remove various components from ProCoder, including the compiler, SQL retriever, and semantic retriever. Notably, the configuration that excludes the compiler and SQL, relying solely on semantic retrieval, is known as RepoCoder Zhang et al. (2023a).
Tables A4, A5, and A6 demonstrate the Pass Rates of ProCoder across CoderEval-python. The data indicates that incorporating Compiler Feedback does improve accuracy, although not markedly substantial. The reason is that highlighting compilation errors alone does not provide related context for resolving them, which is important in project-level coding problems. Also, compared to RepoCoder—which relies solely on Semantic Retrieval, integrating Compiler Feedback with contextual information results in a dramatic performance improvement.
D.2 Varying the Number of Iterations
Table A7: Pass rates of ProCoder on varying iterations.
Iteration Class Level File Level Project Level
$i=0$ 34.55 30.88 13.04
$i=3$ 49.09 47.06 39.13
$i=10$ 47.27 45.59 39.13
To investigate whether iteration could continuously improve performance, we increased the number of iterations to ten. As shown in Table A7, with more iterations, the performance of ProCoder reaches a plateau. This indicates that there are still errors that cannot got repaired by ProCoder, leaving for further research.
D.3 ProCoder on Function-level Code Generation
Table A8: The performance of CoCoGen across varying iterations on the HumanEval benchmark. (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 61.52 80.56 83.15 40 2.38%
CF $i=1$ 71.46 82.46 85.2 7 0.42%
CF $i=2$ 71.4 82.56 85.98 3 0.18%
CF $i=3$ 71.65 82.49 85.36 3 0.18%
Table A9: The performance of CoCoGen across varying iterations on the MBPP benchmark. (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 49.71 58.50 60.81 185 1.90%
CF $i=1$ 52.95 59.58 61.85 11 0.11%
CF $i=2$ 53.06 59.87 62.01 9 0.09%
CF $i=3$ 52.79 59.80 62.09 7 0.07%
To assess the efficacy of ProCoder in function-level code generation, we evaluated it on two commonly used datasets: HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). We retained the compiler feedback model in ProCoder, and conduct tests in a 0-shot scenario, generating 10 code samples for each of the 163 test cases for HumanEval and 974 cases for MBPP, totaling 11,370 samples. The generation and refinement loop iterates 3 times. As shown in Table A8 and Table A9.
The data presented in the table illustrates that compilation feedback improves the accuracy of the generated code, although its efficacy is somewhat limited. We manually review 7 compilation errors not resolved in HumanEval after the first iteration, and find that 6 out of 7 instances were due to LLM erroneously generating code segments enclosed in Markdown code block markers (“‘), resulting in compilation failures. Additionally, another instance of compilation error was identified in the following generated solution:
This code snippet contains five left parentheses and four right parentheses in the return statement, causing the syntax error. Interestingly, despite the compiler has indicated this syntax error, it was not rectified over three iterations.
In MBPP, we manually observe the 7 errors not fixed after the last iteration, and identifies that two test tasks were responsible for all 7 errors: one named "sum" contributed to six errors, and another called "month_season" resulted in one error. We detail these error cases as follows:
D.4 Efficiency of Static Analysis
@@
Utilizing static analysis tools to inspect code typically results in increased latency. To investigate whether the latency impacts the usability of ProCoder, we measure its latency at each test case. We have observed that in experiments, the speed of static analysis and structural query is relatively fast. It is because the semantic checker (i.e., pylint), only analyzes the current file and its dependent files. The average latency was recorded at 1.27 seconds, with a minimum of 0.359 seconds and a maximum of 6.984 seconds. We deem this latency to be acceptable.
Appendix E Algorithm for Project Database Construction
We present the comprehensive algorithm for constructing the project database and generating code with ProCoder.
@@ Appendix E Algorithm for Project Database Construction
The algorithm for building the project database is detailed in Algorithm 1. It involves identifying source files in the project by extracting all files that end with a .py extension. To parse Python source files, we employ the tree-sitter-python parser for generating abstract syntax trees and codeql-python to extract the property of a context entry node. For passage encoding, we utilize text-embedding-ada-002, a model provided by OpenAI and accessible via an online API.

### v3

@@ Appendix C More Details on the Evaluation Metrics
Table 4: A complete list of frequently occurred errors reported by the compiler
@@
Appendix D Errors Reported by the Compiler
@@
Each error is identified by an error code and described with an error message. The error code is a numerical identifier that reflects the error’s nature, while the error message provides a concise description of the error. pylint recognizes 133 distinct error types, and only 12 of these are reported frequently on the CoderEval-Python dataset. We have categorized these errors into four groups based on their characteristics, detailed in Table 1. The FUNC error category is also included in the error distribution analysis but is excluded from the compiler feedback pipeline. This exclusion is because such errors can only be identified with execution, a process that extends beyond the scope of static analysis and is unsuitable for the real-time generation and refinement pipeline. The specific types of each error and corresponding error codes are documented in Table 4.
Table 5: Pass rates on CoderEval-Python of each approach with or without compiler feedback
Method Pass@1 Pass@5 Pass@10
Direct 20.65 26.66 29.13
Direct+CF 32.34 38.43 40.43
ReACC 34.13 41.44 43.48
ReACC+CF 36.60 44.05 46.52
RepoCoder 36.82 40.73 42.17
RepoCoder+CF 38.00 45.38 48.26
Table 6: Pass rates of CoCoGen on Class Runnable test cases
Data Split Class Runnable
Method Pass@1 Pass@5 Pass@10
CoCoGen 28.00 44.92 49.09
- w/o CF and SQL (RepoCoder) 35.45 40.46 41.82
- w/ CF, w/o SQL, w/o Semantic 30.00 42.58 45.45
- w/ CF and SQL, w/o Semantic 30.36 44.61 49.09
- w/ CF and Semantic, w/o SQL 31.45 45.00 47.27
Table 7: Pass rates of CoCoGen on File Runnable test cases
Data Split File Runnable
Method Pass@1 Pass@5 Pass@10
CoCoGen 30.29 43.58 47.06
- w/o CF and SQL (RepoCoder) 29.41 34.61 36.76
- w/ CF, w/o SQL, w/o Semantic 26.03 37.41 42.65
- w/ CF and SQL, w/o Semantic 26.76 37.68 39.71
- w/ CF and Semantic, w/o SQL 27.35 39.90 45.45
Table 8: Pass rates of CoCoGen on Project Runnable test cases
Data Split Project Runnable
Method Pass@1 Pass@5 Pass@10
CoCoGen 21.30 36.73 39.13
- w/o CF and SQL (RepoCoder) 16.96 19.57 21.74
- w/ CF, w/o SQL, w/o Semantic 14.35 25.62 30.43
- w/ CF and SQL, w/o Semantic 15.65 25.12 30.43
- w/ CF and Semantic, w/o SQL 17.83 28.50 34.78
Figure 8: An error case of degenrate solution
Figure 9: An error case of misinterpreting task requirements
Appendix E More Details About the Error Cases
We delve into additional cases where CoCoGen struggles to generate viable solutions, elucidating the underlying causes to foster further research. We identify two primary factors leading to execution errors despite successful compilation. The first factor is the occurrence of degenerate solutions, such as overly simplistic or redundant code. This phenomenon is discovered and explored in Zhang et al. (2023c). Figure 8 exemplifies this issue, showcasing a solution that omits necessary validity checks and error handling, neither of which are specified in the task requirements, nor produced by CoCoGen’s outputs. Our analysis reveals that degenerate solutions frequently pass compilation, rendering compiler-based verification ineffective.
Another prevalent mistake involves misinterpreting task requirements, resulting in solutions that lack logical coherence. Figure 9 depicts an instance where the assignment is to leverage the pre-existing Point class; instead, the LM disregards this specification and redundantly recreates the class. This underscores the challenge LLMs face in accurately comprehending prompts and generating appropriate solutions.
Appendix F More Analytic Experiments
F.1 Usefulness of Compiler Feedback
Here, we examine the usefulness of compiler feedback by integrating it into three baseline models: Vanilla, ReACC, and RepoCoder, each considered separately. We conduct experiments using the GPT-3.5-Turbo and subsequently reported the average score across the entire CoderEval-Python dataset consisting of six levels of context dependencies, as shown in Table 5. The table provides clear evidence that incorporating compiler feedback yields a notable enhancement in model performance for code generation. Specifically, a comparison between RepoCoder with and without compiler feedback reveals a substantial increase in Pass@$1$ from 36.82 to 38.00.
F.2 More Results on the Ablation Study
To further examine the performance of each component of CoCoGen across different dependency levels of test cases, we evaluate CoCoGen at various dependency levels from the CoderEval-Python test suite. The results are presented in Table 6, Table 7, and Table 8. The results show that utilizing compiler feedback and structural queries consistently demonstrates performance improvements across most scenarios. Furthermore, it achieves better performance when stronger cross-file dependencies are present (i.e., at the file-level and project-level tasks), indicating that CoCoGen can accurately capture the context across the project.
F.3 Varying the Number of Iterations
Table 9: Pass rates of CoCoGen on varying iterations
Iteration Class Level File Level Project Level
$i=0$ 34.55 30.88 13.04
$i=3$ 49.09 47.06 39.13
$i=10$ 47.27 45.59 39.13
To investigate whether iteration could continuously improve performance, we increase the number of iterations to ten. As shown in Table 9, with more iterations, the performance of CoCoGen reaches a plateau. This indicates that there are still errors that cannot got repaired by CoCoGen, leaving for further research.
F.4 CoCoGen on Project-level Code Completion
We also evaluate CoCoGen on project-level code completion tasks. Although CoCoGen is originally designed for code generation tasks, the compiler feedback may also benefit this code completion task. The results are presented in Table 10. From the table, we can see that the compiler feedback does improve performance, although not as significantly as it does in the code generation task.
Table 10: C-EM, C-ES, I-EM, and I-F1 scores based on the GPT-3.5-Turbo on the CrossCodeEval dataset
Category Code Match Identifier Match
Method C-EM C-ES I-EM I-F1
Direct 1.91 50.51 3.60 32.27
ReAcc 5.48 54.03 10.02 38.49
RepoCoder 8.52 55.05 13.02 41.05
CoCoGen 9.08 55.31 14.11 42.34
F.5 CoCoGen on Function-level Code Generation
Table 11: The performance of CoCoGen across varying iterations on the HumanEval benchmark (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 61.52 80.56 83.15 40 2.38%
CF $i=1$ 71.46 82.46 85.20 7 0.42%
CF $i=2$ 71.40 82.56 85.98 3 0.18%
CF $i=3$ 71.65 82.49 85.36 3 0.18%
Table 12: The performance of CoCoGen across varying iterations on the MBPP benchmark (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 49.71 58.50 60.81 185 1.90%
CF $i=1$ 52.95 59.58 61.85 11 0.11%
CF $i=2$ 53.06 59.87 62.01 9 0.09%
CF $i=3$ 52.79 59.80 62.09 7 0.07%
To assess the efficacy of CoCoGen in function-level code generation, we evaluate it on two commonly used datasets: HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). We retain the compiler feedback model in CoCoGen, and conduct tests in a 0-shot scenario, generating 10 code samples for each of the 163 test cases for HumanEval and 974 cases for MBPP, totaling 11,370 samples. The generation and refinement loop iterates 3 times, as shown in Table 11 and Table 12.
The data presented in the table illustrates that compilation feedback improves the accuracy of the generated code, although its efficacy is somewhat limited. We manually review 7 compilation errors not resolved in HumanEval after the first iteration, and find that 6 out of 7 instances are caused by the LLM mistakenly generating code segments enclosed in Markdown code block markers (‘‘‘), resulting in compilation failures. Additionally, another instance of compilation error is identified in the following generated solution:
This code snippet contains five left parentheses and four right parentheses in the return statement, causing the syntax error. Interestingly, despite the compiler has indicated this syntax error, it is not rectified over three iterations.
In MBPP, we manually observe the 7 errors not fixed after the last iteration, and identify that two test tasks are responsible for all 7 errors: one named "sum" contributed to six errors, and another called "month_season" resulted in one error. We detail these error cases as follows:
Table 13: Tables pre-computed by CoCoGen for error correction
@@ F.5 CoCoGen on Function-level Code Generation
Table 14: A complete list of demonstration examples prompted to the language model
@@
F.6 Efficiency of Static Analysis
@@
Utilizing static analysis tools to inspect code typically results in increased latency. To investigate whether the latency impacts the usability of CoCoGen, we measure its latency at each test case. We have observed that in experiments, the speed of static analysis and structural query is relatively fast. It is because the semantic checker (i.e., pylint), only analyzes the current file and its dependent files. The average latency is recorded at 1.27 seconds, with a minimum of 0.359 seconds and a maximum of 6.984 seconds. We deem this latency to be acceptable.
Appendix G Technical Details of Structural Query System
G.1 The Design Principle
In CoCoGen, we employ the Structural Query Language (SQL) to perform structured queries on code repositories. The SQL syntax used here is CodeQL, a specialized version designed for software repository mining, capable of conducting complex control-flow and data-flow analyses on a set of code files.
We extracted 11 error codes from the compiler (shown in Table 4), each corresponding to one of four error categories as in Section 2.2. For efficiency, two authors manually inspected the most frequently occurring error codes, precomputed six data tables from the project graph, and hard-coded the structural context retrieval procedure for the following error codes: E0001 (syntax error), E0602 (undefined-variable), E1101 (no-member), E0213 (no-self-argument), and E0102 (function redefined). The tables to retrieve structural project context for these frequently occurred errors are presented in Table 13.
Whenever the compiler reports an error, if it is a frequently occurring error, its retrieval is done from these precomputed data tables. Otherwise, the CoCoGen invokes the LLM to generate an SQL query statement, which is then executed on the project graph.
In the CoderEval-Python benchmark, with iteration rounds set to 3, related project contexts of 93.2% of errors are successfully retrieved. For the remaining 6.8% of cases, there are two reasons for failure: first, the LLM does not follow the demonstrated CodeQL examples and generates queries that are syntactically or semantically incorrect, thus rejected by the CodeQL query system; second, the query does not return anything because the specified conditions in the query are not satisfied in project context.
G.2 Demonstration Examples of Structural Queries
We present a demonstration example in Section 3.3 to compose the structural query. The example focuses on identifying and addressing instances of missing or incorrectly utilized context entries. We detail several example error messages along with their corresponding structural queries. A comprehensive list of these examples can be found in Table 14. The four queries shown in the table are written manually by one of the authors, and designed to handle four representative types of compilation errors. These sampled error messages are reported by compilers when generating solutions without project context. The specific four messages are chosen randomly from the compilation log.
Appendix H Algorithm for Project Database Construction
@@
We present the comprehensive algorithm for constructing the project database and generating code with CoCoGen. The algorithm for building the project database is detailed in Algorithm 1. It involves identifying source files in the project by extracting all files that end with a .py extension. To parse Python source files, we employ the tree-sitter-python parser for generating abstract syntax trees and codeql-python to extract the property of a context entry node. To encode passages to vectors, we utilize text-embedding-ada-002, a text embedding model provided by OpenAI and accessible via online APIs OpenAI (2022).